# SALAD: Self-Adaptive Link Adaptation

### v1

@@ Abstract
Adapting the modulation and coding scheme (MCS) to the wireless link quality is critical for maximizing spectral efficiency while ensuring reliability. We propose SALAD (self-adaptive link adaptation), an algorithm that exclusively leverages ACK/NACK feedback to reliably track the evolution of the signal-to-interference-plus-noise ratio (SINR), achieving high spectral efficiency while keeping the long-term block error rate (BLER) near a desired target. SALAD infers the SINR by minimizing the cross-entropy loss between received ACK/NACKs and predicted BLER values, with a learning rate that self-adapts online through knowledge distillation. Based on this inference, SALAD selects the MCS via hypothesis testing: if the SINR is likely underestimated, a higher MCS is selected to accelerate link adaptation under improving channel conditions. To prevent BLER drift from its long-term target, SALAD incorporates a feedback control loop that adjusts the instantaneous BLER target. Over-the-air experiments on a 5G testbed demonstrate that SALAD consistently outperforms the industry-standard outer-loop link adaptation (OLLA). With a single set of parameters, SALAD achieves up to 15% higher throughput and spectral efficiency than multiple OLLA variants across different traffic regimes, while meeting the BLER target.
@@ I-B Our contribution
We propose self-adaptive link adaptation (SALAD), a novel algorithm that estimates the SINR by minimizing the cross-entropy loss between the observed ACK/NACK feedback and the BLER via gradient descent, as described in Section III-A. Its learning rate can be adjusted in an online fashion by a mechanism inspired by knowledge distillation, which compares the gradient-based estimate to a more powerful batch estimator. When the SINR is deemed to be underestimated—based on a hypothesis testing—SALAD increases the instantaneous BLER target to probe a higher MCS value and corrects the SINR estimation accordingly (Section III-B). To meet a desired long-term BLER, SALAD adjusts the instantaneous BLER target via a feedback loop (Section III-C).
@@
The code used to produce the results presented in Section IV-C is publicly available at [9].
@@
$\text{\small{BLER}}^{\mathrm{tab}}$ Block error rate (BLER) from pre-computed table
@@
$\varepsilon$ SALAD’s student model learning rate
@@ II The industry standard: OLLA
$\operatorname{argmax}\big\{u\!:\text{\small{BLER}}^{\mathrm{tab}}\!\left(u,\widetilde{\gamma}^{\mathrm{OLLA}}_{t-\delta_{t}},b_{t}\right)\leq\tau\big\},$ (3)
where $\text{\small{BLER}}^{\mathrm{tab}}$ contains pre-computed BLER values, simulated on the AWGN channel. This step is commonly called inner-loop link adaptation (ILLA) [^4] An alternative, drop-in replacement for ILLA selects the MCS that maximizes the expected spectral efficiency in (3)..
@@ III-A SINR inference
We first illustrate how SALAD produces an estimate $\widetilde{\gamma}_{t}$ of the SINR at slot $t$ from past MCS values and the corresponding ACK/NACK feedback. Our method works across two different time-scales. At every slot, a simple “student” iterate infers the next SINR from the last estimate. Such iterate depends on a learning rate called $\varepsilon$. Optionally, every $N_{\varepsilon}$ slots, a powerful “teacher” model infers the whole time-series of past SINR. The student’s learning rate is finally optimized to best mimic the teacher’s behavior over the past slots. This allows SALAD to capture long-term SINR volatility patterns. We start by presenting our teacher model.
Teacher model. We introduce a parametric model $\widetilde{\gamma}_{i}(\theta)$ of the estimated SINR at each past slot $i\leq t$, parametrized by $\theta\in\mathbb{R}^{K}$. To infer the SINR evolution, we seek the parameters $\theta$ that minimize the discrepancy between the observed ACK/NACKs and the corresponding likelihood, quantified by the empirical binary cross-entropy (BCE) loss. Our “teacher” SINR inference formulation is as follows:
$\theta^{*}_{t}:=\operatorname{argmin}_{\theta}\ \sum_{i<t}\text{\small{BCE}}_{i}(\theta)+\Omega(\theta),$ (8)
where $\Omega(.)$ is a regularizer, promoting temporal smoothness of the estimated SINR, and
$\text{\small{BCE}}_{i}(\theta):=$ $\,-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}\log\Pr\left(1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}=1\,|\,u_{i},\widetilde{\gamma}_{i}(\theta),b_{i}\right)$
$-\big(1$- $1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}\big)\log\Pr\left(1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}=0\,|\,u_{i},\widetilde{\gamma}_{i}(\theta),b_{i}\right).$ (9)
The probability of receiving a NACK in an equivalent AWGN channel can be pre-computed and stored in BLER tables $\text{\small{BLER}}^{\mathrm{tab}}$, for every MCS $u$, SINR $\gamma$ and TBS $b$. Then, (9) can be rewritten as
$\text{\small{BCE}}_{i}(\theta)=$ $\,-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}\log\text{\small{BLER}}^{\mathrm{tab}}_{i}(\theta)$
$-\big(1-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}\big)\log\left(1-\text{\small{BLER}}^{\mathrm{tab}}_{i}(\theta)\right),$ (10)
where we define:
$\text{\small{BLER}}^{\mathrm{tab}}_{i}(\theta):=\text{\small{BLER}}^{\mathrm{tab}}\left(u_{i},\widetilde{\gamma}_{i}(\theta),b_{i}\right).$ (11)
Gradient computation: To solve (8) via gradient descent, we must compute the derivative of the loss with respect to $\theta$. To obtain a convenient closed form, we approximate $1-\text{\small{BLER}}^{\mathrm{tab}}$ as a suitably translated and scaled sigmoid $\sigma$, as in [30]:
$1\!-\!\text{\small{BLER}}^{\mathrm{tab}}(u,\gamma,b)\!\approx\!\left(1+e^{-\frac{\gamma-c(u,b)}{s(u,b)}}\right)^{-1}\!\!:=\!\sigma_{u,b}(\gamma),$ (12)
@@
By exploiting the property $\sigma^{\prime}=\frac{1}{s}\sigma(1-\sigma)$, we obtain
$\nabla_{\theta}\text{\small{BCE}}_{i}(\theta)\approx s_{i}^{-1}\left(1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}-\text{\small{BLER}}^{\mathrm{tab}}_{i}(\theta)\right)\nabla_{\theta}\widetilde{\gamma}_{i}(\theta).$ (13)
Piece-wise linear model: A simple yet effective model is the piece-wise linear one, also called first-order spline:
$\widetilde{\gamma}_{i}(\theta)=\sum_{k=1}^{K}\widetilde{\gamma}^{k}(i,\theta_{k}),$ (14)
where $\widetilde{\gamma}^{k}(i,\theta_{k})$ is a triangle-shaped function centered at $t_{k}$ with peak height $\theta_{k}$:
$\widetilde{\gamma}^{k}(i,\theta_{k}):=\left\{\begin{array}[]{l}\theta_{k}\tfrac{t_{k+1}-i}{t_{k+1}-t_{k}}\quad\text{if }i\in[t_{k},t_{k+1}]\\ \theta_{k}\frac{i-t_{k-1}}{t_{k}-t_{k-1}}\quad\text{if }i\in[t_{k-1},t_{k}]\\ 0\quad\text{otherwise.}\end{array}\right$. (15)
In this case, a reasonable regularizer is the total variation one:
$\Omega(\theta):=\beta\sum_{k=2}^{K}\left(\theta_{k}-\theta_{k-1}\right)^{2},\quad\beta\geq 0.$ (16)
For the linear model, the gradients $\nabla_{\theta}\widetilde{\gamma}_{i}(\theta)$, $\nabla_{\theta}\Omega(\theta)$ write:
$\tfrac{\partial}{\partial\theta_{k}}\gamma_{i}(\theta)=$ $\,\left\{\begin{array}[]{l}\tfrac{t_{k+1}-i}{t_{k+1}-t_{k}}\quad\text{if }i\in[t_{k},t_{k+1}]\\ \frac{i-t_{k-1}}{t_{k}-t_{k-1}}\quad\text{if }i\in[t_{k-1},t_{k}]\\ 0\qquad\quad\ \ \text{otherwise}\end{array}\right.,\ \forall k$ (20)
$\tfrac{\partial}{\partial\theta_{k}}\Omega(\theta)=$ $\left\{\begin{array}[]{l}2\beta(2\theta_{k}-\theta_{k-1}-\theta_{k+1}),\ 1<k<K\\ 2\beta(\theta_{k}-\theta_{k+1}),\qquad\qquad k=1\\ 2\beta(\theta_{k}-\theta_{k-1}),\qquad\qquad k=K.\end{array}\right$. (24)
The values of $\alpha$ and $K$ can be determined by cross-validation.
The time-series $\{\widetilde{\gamma}_{i}(\theta^{*}_{t})\}_{i<t}$, where $\theta^{*}_{t}$ is the solution to (8), is our teacher model. The SINR estimation at slot $t$ can be extrapolated as $\widetilde{\gamma}_{t}(\theta^{*}_{t})$.
@@
Student model. It may be computationally prohibitive to re-optimize the teacher model at every slot to produce the estimate $\widetilde{\gamma}_{t}(\theta^{*}_{t})$. We then propose a computationally cheaper but effective heuristic alternative, called student inference, which infers the SINR at slot $t$ as a gradient update of the previous estimate:
@@
where $\varepsilon>0$ is the learning rate and $\widetilde{\gamma}_{0}$ is initialized arbitrarily. By using the sigmoid approximation (12) and exploiting the property $\sigma^{\prime}=\frac{1}{s}\sigma(1-\sigma)$, we can rewrite (25) in closed-form:
$\widetilde{\gamma}_{t}=\widetilde{\gamma}_{t-1}+\varepsilon s_{t-1}^{-1}\left(\text{\small{BLER}}^{\mathrm{tab}}_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}\right)\quad\forall\,t>0,$ (26)
where we define:
$\text{\small{BLER}}^{\mathrm{tab}}_{t}:=\text{\small{BLER}}^{\mathrm{tab}}\left(u_{t},\widetilde{\gamma}_{t},b_{t}\right)\quad\forall\,t\geq 0$ (27)
and $s_{t}:=s(u_{t},b_{t})$ is the sigmoid scale corresponding to MCS $u_{t}$ and TBS $b_{t}$.
Surprise effect: We observe that the student’s SINR increment in (26) is proportional to the term $\text{\small{BLER}}^{\mathrm{tab}}_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}$ which quantifies the “surprise” of observing $1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}$ relative to its expected likelihood $\text{\small{BLER}}^{\mathrm{tab}}_{t-1}$. Indeed,
  • Suppose $\text{\small{BLER}}^{\mathrm{tab}}_{t-1}\approx 1$:
@@
  • Conversely, if $\text{\small{BLER}}^{\mathrm{tab}}_{t-1}\approx 0$:
@@
Knowledge distillation. For the student SINR estimate (26) to accurately track actual SINR variations, the learning rate $\varepsilon$ should ideally increase with the volatility of the SINR. Yet, the ground truth is inherently unobservable. To overcome this, we adopt a procedure inspired by knowledge distillation in neural networks [32], where the student estimator learns to imitate a more powerful teacher estimator. Specifically, the teacher SINR estimator is first optimized by solving (8) according to recent observations every $N_{\varepsilon}$ slots. Then, the student estimator (26) is replayed on the same data using different values of $\varepsilon$, and the value that yields the closest match to the reference estimator is selected for use in the subsequent slots. This procedure is described in Algorithm 2.
Algorithm 2 Knowledge Distillation to Adapt Learning Rate
1:At slot $t$, optimize teacher model (8). Obtain $\widetilde{\gamma}_{i}(\theta^{*}),\,\forall i\!\leq\!t$
2:Replay (26), for different $\varepsilon$; obtain $\widetilde{\gamma}^{(\varepsilon)}_{i},\ \forall\,i\leq t$ and $\forall\,\epsilon$
3:Select student’s learning rate as
$\varepsilon^{*}=\operatorname{argmin}_{\varepsilon}\sum_{i<t}\left(\widetilde{\gamma}_{i}(\theta^{*})-\widetilde{\gamma}_{i}^{(\varepsilon)}\right)^{2}.$
@@ III-B MCS selection
  • If the SINR is detected to be underestimated, a high instantaneous BLER target $\tau_{t}:=\tau^{\mathrm{probe}}\gg\tau$ is probed. This corresponds to a high MCS index and a high BLER ($\text{\small{BLER}}^{\mathrm{tab}}_{t}\approx 1$). Receiving an ACK ($1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t}=0$) in this setting provides two advantages:
@@
$\mathcal{S}_{t}=\frac{1}{T}\sum_{i=1}^{T}\text{\small{BLER}}^{\mathrm{tab}}_{t-\delta_{t}-i}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-\delta_{t}-i}.$ (29)
@@
$\mathcal{H}_{0,t}:=\!\left\{\mathbb{E}[1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i-\delta_{t}}]\!=\!\text{\small{BLER}}^{\mathrm{tab}}_{i-\delta_{t}},\ t-T\leq i\leq t-1\right\}.$ (30)
@@
$\mathrm{Var}\left[\mathcal{S}_{t}|\mathcal{H}_{0,t}\right]=\frac{1}{T^{2}}\sum_{i=1}^{T}\text{\small{BLER}}^{\mathrm{tab}}_{t-\delta_{t}-i}(1-\text{\small{BLER}}^{\mathrm{tab}}_{t-\delta_{t}-i}).$ (32)
@@ III-C Long-term BLER target enforcement
8:# Student’s SINR inference update
@@
$\widetilde{\gamma}\leftarrow\widetilde{\gamma}+\varepsilon s^{-1}\left(\text{\small{BLER}}^{\mathrm{tab}}(u,\widetilde{\gamma},b)-1{\thinspace}\mathrm{I}^{\textsc{NACK}}\right)$
@@
26:  if $\mod(t,N_{\varepsilon})=0$ then
27:    # Knowledge distillation
28:   Update the learning rate $\varepsilon$ via Algorithm 2
29:  end if
30:end for
31:Return $u_{t}$
@@ III-D A practical consideration
As a result, the term $s_{t-1}^{-1}\left(\text{\small{BLER}}^{\mathrm{tab}}_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}\right)$, which appears in SALAD’s SINR update (26), can take on extreme values when the TBS is large. Specifically, upon probing, a received ACK leads to a substantial SINR increase $\widetilde{\gamma}_{t}\gg\widetilde{\gamma}_{t-1}$. Conversely, receiving a NACK results in only a negligible update $\widetilde{\gamma}_{t+1}\approx\widetilde{\gamma}_{t}$. In the absence of probing, the situation is reversed: an ACK leads to a negligible SINR update, while a NACK causes a significant SINR decrease.
A simple and effective remedy is to clip the values of $\text{\small{BLER}}^{\mathrm{tab}}_{t}$ and $s_{t}$ to appropriate intervals, e.g., $(0.01,0.99)$ and $(0.5,10)$, respectively.
@@ III-E The impact of delayed HARQ feedback
$\tau_{t-1}\approx\text{\small{BLER}}^{\mathrm{tab}}_{t-1}:=\text{\small{BLER}}^{\mathrm{tab}}(u_{t-1},\widetilde{\gamma}_{t-1}),\ \text{if }\delta=0.$ (35)
@@
Presence of HARQ feedback delay ($\delta>0$). In this scenario, SALAD’s SINR update differs significantly from that of time-adaptive OLLA. In both algorithms, the MCS $u_{t-1}$ is selected based on an outdated SINR estimation, specifically, $u_{t-1}=u^{\mathrm{ILLA}}_{t-1}(\widetilde{\gamma}_{t-1-\delta_{t-1}},\tau_{t-1})$. Yet, the SINR staleness renders (35) invalid: the target $\tau_{t-1}$, used by OLLA in its SINR update, is different from the effective BLER, $\text{\small{BLER}}^{\mathrm{tab}}_{t-1}$, used by SALAD. In other words, SALAD updates the SINR using fresher information than OLLA. As a result, SALAD’s SINR estimation is more robust to feedback delay, as confirmed by the experiments in Section IV-A.

### v2

@@ Abstract
Adapting the modulation and coding scheme (MCS) to the wireless link quality is critical for maximizing spectral efficiency while ensuring reliability. We propose SALAD (self-adaptive link adaptation), an algorithm that exclusively leverages ACK/NACK feedback to reliably track the evolution of the signal-to-interference-plus-noise ratio (SINR), achieving high spectral efficiency while keeping the long-term block error rate (BLER) near a desired target. SALAD infers the SINR by minimizing the cross-entropy loss between received ACK/NACKs and predicted BLER values. Based on this inference, SALAD selects the MCS via hypothesis testing: if the SINR is likely underestimated, a higher MCS is selected to accelerate link adaptation under improving channel conditions. To prevent BLER drift from its long-term target, SALAD incorporates a feedback control loop that adjusts the instantaneous BLER target. Over-the-air experiments on a 5G testbed demonstrate that SALAD consistently outperforms the industry-standard outer-loop link adaptation (OLLA). With a single set of parameters, SALAD achieves up to 15% higher throughput and spectral efficiency than multiple OLLA variants across different traffic regimes, while meeting the BLER target.
@@ I-B Our contribution
We propose self-adaptive link adaptation (SALAD), a novel algorithm that estimates the SINR by minimizing the cross-entropy loss between the observed ACK/NACK feedback and the BLER via gradient descent, as described in Section III-A. When the SINR is deemed to be underestimated—based on a hypothesis testing—SALAD increases the instantaneous BLER target to probe a higher MCS value and corrects the SINR estimation accordingly (Section III-B). To meet a desired long-term BLER, SALAD adjusts the instantaneous BLER target via a feedback loop (Section III-C).
@@
An implementation of SALAD is publicly available at [19].
@@
BLER Block error rate (BLER) from pre-computed table
@@
$\varepsilon$ SALAD’s learning rate
@@ II The industry standard: OLLA
$\operatorname{argmax}\big\{u\!:\text{\small{BLER}}\!\left(u,\widetilde{\gamma}^{\mathrm{OLLA}}_{t-\delta_{t}},b_{t}\right)\leq\tau\big\},$ (3)
where BLER contains pre-computed BLER values, simulated on the AWGN channel. This step is commonly called inner-loop link adaptation (ILLA) [^4] An alternative, drop-in replacement for ILLA selects the MCS that maximizes the expected spectral efficiency in (3)..
@@ III-A SINR inference
We illustrate how SALAD produces an estimate $\widetilde{\gamma}_{t}$ of the SINR at slot $t$ from past MCS values and the corresponding ACK/NACK feedback.
First, we quantify the discrepancy between the observed ACK/NACKs in slot $i$ and the SINR estimate $\widetilde{\gamma}_{i}$ via the empirical binary cross-entropy (BCE) loss, defined as
$\text{\small{BCE}}_{i}(\widetilde{\gamma}_{i}):=$ $\,-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}\log\text{\small{BLER}}\left(u_{i},\widetilde{\gamma}_{i},b_{i}\right)$
$-\big(1$- $1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i}\big)\log\text{\small{BLER}}\left(u_{i},\widetilde{\gamma}_{i},b_{i}\right).$ (8)
We seek to minimize the sum of the BCE loss terms across slots:
$\min_{\{\widetilde{\gamma}\}_{i\geq 1}}\ \sum_{i\geq 1}\text{\small{BCE}}_{i}(\widetilde{\gamma}_{i}).$ (9)
We need to solve (9) in an online setting: to estimate the SINR for the next slot, we can only rely on past ACK/NACK feedback. To this aim, we shall resort to the framework of online optimization. To obtain a closed-form expression of the SINR update, we conveniently approximate $1-\text{\small{BLER}}$ as a suitably translated and scaled sigmoid $\sigma$, as in [7]:
$1\!-\!\text{\small{BLER}}(u,\gamma,b)\!\approx\!\left(1+e^{-\frac{\gamma-c(u,b)}{s(u,b)}}\right)^{-1}\!\!:=\!\sigma_{u,b}(\gamma),$ (10)
@@
Under approximation (10), the BCE loss is convex. We can then apply the online convex optimization paradigm [13]. Specifically, we infer the SINR at slot $t$ via online gradient descent as:
@@
where $\varepsilon>0$ is the learning rate and $\widetilde{\gamma}_{0}$ is initialized arbitrarily. We can further express the loss gradient analytically and rewrite (11) in closed-form:
$\widetilde{\gamma}_{t}=\widetilde{\gamma}_{t-1}+\varepsilon s_{t-1}^{-1}\left(\text{\small{BLER}}_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}\right)\quad\forall\,t>0,$ (12)
where $\text{\small{BLER}}_{t}:=\text{\small{BLER}}\left(u_{t},\widetilde{\gamma}_{t},b_{t}\right)$ and $s_{t}:=s(u_{t},b_{t})$.
Surprise effect: We observe that the SINR increment in (12) is proportional to the term $\text{\small{BLER}}_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}$ which quantifies the “surprise” of observing $1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}$ relative to its expected likelihood $\text{\small{BLER}}_{t-1}$. Indeed,
  • Suppose $\text{\small{BLER}}_{t-1}\approx 1$:
@@
  • Conversely, if $\text{\small{BLER}}_{t-1}\approx 0$:
@@ III-B MCS selection
  • If the SINR is detected to be underestimated, a high instantaneous BLER target $\tau_{t}:=\tau^{\mathrm{probe}}\gg\tau$ is probed. This corresponds to a high MCS index and a high BLER ($\text{\small{BLER}}_{t}\approx 1$). Receiving an ACK ($1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t}=0$) in this setting provides two advantages:
@@
$\mathcal{S}_{t}=\frac{1}{T}\sum_{i=1}^{T}\text{\small{BLER}}_{t-\delta_{t}-i}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-\delta_{t}-i}.$ (14)
@@
$\mathcal{H}_{0,t}:=\!\left\{\mathbb{E}[1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{i-\delta_{t}}]\!=\!\text{\small{BLER}}_{i-\delta_{t}},\ t-T\leq i\leq t-1\right\}.$ (15)
@@
$\mathrm{Var}\left[\mathcal{S}_{t}|\mathcal{H}_{0,t}\right]=\frac{1}{T^{2}}\sum_{i=1}^{T}\text{\small{BLER}}_{t-\delta_{t}-i}(1-\text{\small{BLER}}_{t-\delta_{t}-i}).$ (17)
@@ III-C Long-term BLER target enforcement
8:# SINR inference update
@@
$\widetilde{\gamma}\leftarrow\widetilde{\gamma}+\varepsilon s^{-1}\left(\text{\small{BLER}}(u,\widetilde{\gamma},b)-1{\thinspace}\mathrm{I}^{\textsc{NACK}}\right)$
@@
26:end for
27:Return $u_{t}$
@@ III-D A practical consideration
As a result, the term $s_{t-1}^{-1}\left(\text{\small{BLER}}_{t-1}-1{\thinspace}\mathrm{I}^{\textsc{NACK}}_{t-1}\right)$, which appears in SALAD’s SINR update (12), can take on extreme values when the TBS is large. Specifically, upon probing, a received ACK leads to a substantial SINR increase $\widetilde{\gamma}_{t}\gg\widetilde{\gamma}_{t-1}$. Conversely, receiving a NACK results in only a negligible update $\widetilde{\gamma}_{t+1}\approx\widetilde{\gamma}_{t}$. In the absence of probing, the situation is reversed: an ACK leads to a negligible SINR update, while a NACK causes a significant SINR decrease.
A simple and effective remedy is to clip the values of $\text{\small{BLER}}_{t}$ and $s_{t}$ to appropriate intervals, e.g., $(0.01,0.99)$ and $(0.5,10)$, respectively.
@@ III-E The impact of delayed HARQ feedback
$\tau_{t-1}\approx\text{\small{BLER}}_{t-1}:=\text{\small{BLER}}(u_{t-1},\widetilde{\gamma}_{t-1}),\ \text{if }\delta=0.$ (20)
@@
Presence of HARQ feedback delay ($\delta>0$). In this scenario, SALAD’s SINR update differs significantly from that of time-adaptive OLLA. In both algorithms, the MCS $u_{t-1}$ is selected based on an outdated SINR estimation, specifically, $u_{t-1}=u^{\mathrm{ILLA}}_{t-1}(\widetilde{\gamma}_{t-1-\delta_{t-1}},\tau_{t-1})$. Yet, the SINR staleness renders (20) invalid: the target $\tau_{t-1}$, used by OLLA in its SINR update, is different from the effective BLER, $\text{\small{BLER}}_{t-1}$, used by SALAD. In other words, SALAD updates the SINR using fresher information than OLLA. As a result, SALAD’s SINR estimation is more robust to feedback delay, as confirmed by the experiments in Section IV-A.